How would I write a formula for this and how would I complete it?
"A rectangle is to be drawn with perimeter 64 cm. If the length is to be 14 cm more than the width, determine the area of the rectangle."
Help Please?

Answer by checkley77(12569)

L=W+14
2L+2W=64
2(W+14)=2W=64
2W+28+2W=64
4W=64-28
4W=36
W=35/4
W=9 cm.IS THE WIDTH
L=9+14=23 cm.IS THE LENGTH.
PROOF:
2*9+2*23=64
18+46=64
64=64
L*W=AREA
9*23=207 cm^2 IS THE AREA.